# A Modern Take on Visual Relationship Reasoning for Grasp Planning

### v1

@@ Abstract
Interacting with real-world cluttered scenes pose several challenges to robotic agents that need to understand complex spatial dependencies among the observed objects to determine optimal pick sequences or efficient object retrieval strategies. Existing solutions typically manage simplified scenarios and focus on predicting pairwise object relationships following an initial object detection phase, but often overlook the global context or struggle with handling redundant and missing object relations. In this work, we present a modern take on visual relational reasoning for grasp planning. We introduce D3GD, a novel testbed that includes bin picking scenes with up to 35 objects from 97 distinct categories. Additionally, we propose D3G, a new end-to-end transformer-based dependency graph generation model that simultaneously detects objects and produces an adjacency matrix representing their spatial relationships. Recognizing the limitations of standard metrics, we employ the Average Precision of Relationships for the first time to evaluate model performance, conducting an extensive experimental benchmark. The obtained results establish our approach as the new state-of-the-art for this task, laying the foundation for future research in robotic manipulation. We publicly release the code and dataset at https://paolotron.github.io/d3g.github.io.
@@ I Introduction
Previous research has framed this problem as visual manipulation relationship reasoning [1, 2, 3], highlighting as a primary challenge the inference of relationship among objects, which can be categorized as parent, child, or none. In this context, objects are nodes in a dependency graph, where parent nodes should be grasped after child nodes. Knowing the complete dependency graph enables the calculation of the shortest path to any object and facilitates planning the most efficient picking sequence to access it.
Existing algorithms address this task by either focusing solely on modeling relationships between pairs of objects, thereby neglecting the broader global context, or by inferring the graph from the entire image scene, which often leads to difficulties in handling redundant or missing edges [4, 1, 2]. Additionally, their approach treats object detection and manipulation relationship recognition as two distinct stages, with evaluation primarily concentrated on the latter. The experimental analyses usually focus on per-edge prediction, and typically run on scenes with a low (5 objects, [1]) or medium (19 objects, [5]) number of objects.
@@
  • we introduce a new testbed based on [6]. We organize bin picking scenes in subsets of growing complexity based on the graph that outlines the objects’ dependencies, focusing on the number of nodes, connected subgraphs, and shortest paths from leaf to root. We name it D3GD: Dependency Graph Generation for Grasp planning Dataset, and contains scenes with up to 35 objects belonging to 97 different categories. It includes synthetic images and an additional real-world test split to study sim-to-real transfer;
  • we design a new end-to-end dependency graph generation model. We refer to it as D3G. Its objective is to jointly detect the objects and predict the adjacency matrix describing their spatial relationship. We formalize the task as a set prediction problem over the whole image and learn the edge features representing the graph relationships rather than relying on object localization and their feature post-hoc concatenation as done in previous work;
  • we assess the entire pipeline with a tailored metric. Standard evaluation strategies depend on implicit object scoring thresholds which may be difficult to set. Here we use for the first time the Average Precision of Relationships [7] to evaluate the generated dependency graph.
With a thorough experimental analysis we show how our approach outperforms several state-of-the art competitors in challenging experimental settings. Overall our work presents a novel benchmark suite that will be open-sourced to the community paving the way for future application-oriented research in robotic manipulation.
@@ II Related Works
Object Detection is an extensively studied task whose goal is to localize objects in a scene and recognize their category label. Detection approaches evolved from feature matching and part modeling [8], through two stages [9, 10] and one stage deep methods [11, 12], getting more recently to transformer-based techniques. The latter allows for global computation over a whole image and set prediction, leveraging attention between multiple objects as in the case of DETR [13]. The use of deformable attention has been shown to further boost such approaches [14].
@@
Graph Learning is a powerful tool when studying complex interactions within structured data. Message passing based networks have been designed to leverage the sparsity and connectivity of graph structures, efficiently propagating information and capturing dependencies between nodes [15]. With the advent of self-attention based architectures [16], several efforts have been made to merge the graph neural network paradigm and the transformer one, obtaining competitive performance on graph benchmarks [17, 18, 19].
@@
Relationship reasoning is an umbrella term that encompasses a variety of tasks where it is important to capture structural information from data and to model relations that can be simply geometric (spatial positioning) or semantic (e.g. how the action of a subject affects an object). Large datasets such as Visual Genome [20] and Open Images [7] paved the way for perception studies based on deep learning mainly dedicated to image captioning, with first solutions building on two-stage object detectors [21], and the latest ones based on transformers [22] with promising applications for the robotic field [23, 24].
In the area of decision-making processes and motion planning, generating the dependency graph is a pre-condition to grasping in cluttered scenes. Recent works have also included it as an intermediate [25, 26] or auxiliary objective when predicting robotic parallel grasp poses [27, 28, 29].
@@
The first real-world visual manipulation relationship dataset (VMRD) was introduced in [1]. It is a collection of 5185 hand-annotated images depicting various configurations of objects in cluttered scenes each containing up to 5 object instances sampled from 31 categories. While valid, this dataset is limited in dimension and complexity. The authors also proposed a deep model including distinct stages for object detection and relationship prediction. The same high-level solution was then adopted by following approaches that used fully connected conditional random fields [4] and graph neural networks [2, 3]. These works included analyses on a larger relational grasp dataset named REGRAD [5] automatically collected in a virtual environment. It is composed of images annotated with manipulation relationship between objects in dense clutter and including multi-view perspectives [30]. Nonetheless, existing testbeds tend to focus on relatively simple scenarios and often lack comprehensive information on the experimental settings, making previous results difficult to reproduce [^1] Reproducible results on a dataset require knowledge of the pre-processing pipeline and access to baseline code and checkpoints. Unfortunately, most prior work in visual manipulation relationship reasoning lacks these resources. Specifically, in REGRAD, detailed analysis uncovered issues such as missing boxes, one-pixel masks, and inconsistent object classes. The lack of public code further impedes result replication. We overcome these limitations with our benchmark, re-implementing the competitor methods and releasing the code for fair comparison.. Lately with the release of the NVIDIA Isaac simulator [31] a new generation of highly realistic and accurately annotated datasets have been published [32, 6, 33], which provide the opportunity to scale up relational reasoning related tasks for robotics applications.
@@ III Dataset
We define our experimental testbed D3GD to get closer to real-world conditions where robots are challenged with grasping objects from large containers. We build on MetaGraspNetV2 [6] which is composed of nearly 300K synthetic images and 2.3K real images depicting bin picking scenes composed by up to 35 object instances sampled from 97 classes. The dataset comes with both 2D and 3D modalities as well as various types of grasp and labels. For our use-case we consider only RGB images, the object bounding boxes, the grasping dependency graph and, for the baseline models that use them, the instance segmentation annotations.
We organize the dataset into different difficulty levels on the basis of three key metrics: number of objects, largest subgraph, and longest shortest path from leaf to root. When the number of objects in the scene increases, detecting them and reasoning on their relations may become more challenging due to occlusions. Thus, it is important to consider the scene density besides the object cardinality. We do that by measuring how many objects (nodes) are part of a connected subgraph within the whole dependency graph. Finally, we measure the longest shortest paths from a graph leaf (the first object to be picked) to the root (the last one). This means that we count the minimum number of objects that we need to remove before reaching an object at the bottom of the bin. As the dependency graph of a scene may have multiple leaves and roots, we annotate for each scene the longest of these paths which is a good indicator of how many layers of clutter the bin contains. Note that we are interpreting the dependency graph as a Directed Acyclic Graph (DAG) and we removed all the not-fitting scene samples (only $\sim$60) that, upon visual inspection, were clearly identified as errors produced during the synthetic dataset generation phase.
Overall, our D3GD is organized with difficulty sets as presented in Table I. For our study we disregard the trivial set and focus on the easy, medium, and hard ones (see examples in Figure 2). We split the dataset into train, eval, and test scenes with an 80:10:10 proportion.
TABLE I: Difficulty levels decomposition of our D3GD dataset. Each row describes the concurrent conditions that define a level set. For instance, a sample in the medium set is a scene with 6 or more objects, of which more than 40% are involved in connected subgraphs and showing at least 4 stacked objects.
Set # Objects Largest Longest # Samples
Subgraph Shortest Path
@@ IV Method
In this section we present our new visual relationship reasoning model D3G. It is formalized as a transformer-based graph generation approach that predicts the graph nodes by detecting the objects in the scene and the edges representing the object spatial relationships (presence/absence and direction). We depart from commonly used GNNs architecture and propose a fusion between graph transformers and detection transformers. In this way all the components of the dependency graph can be obtained jointly. In contrast to previous works using two-stage pipelines [1, 2, 3], our end-to-end strategy avoids the use of feature-cropping layers that in high clutter environments tend to generate ambiguous features. Figure 3 presents a schematic illustration of the proposed approach and helps to visualize the model flow explained in the following.
Figure 2: Data samples from the three main difficulty levels of D3GD.
@@ IV-A Detecting Objects as Graph Nodes
Given an input image of dimension $3\times H_{0}\times W_{0}$, we perform object detection as a direct set prediction problem by following DETR [13]. A CNN backbone extracts a compact feature representation of dimensionality $C\times H\times W$ (where $C=2048$, $H={H_{0}}/{32}$, $W={W_{0}}/{32}$) which is rescaled via a $1\times 1$ convolution to reduce the channel dimensionality. The resulting $d\times H\times W$ tensor is interpreted as a list of vectors ($d\times HW$). These are fed to a transformer encoder [16] together with their spatial positional encodings and the obtained output is passed to the transformer decoder which also receives $N$ object queries (tensor of dimensionality $N\times F_{0}$, initially set to zero). The latter are transformed by the decoder in new embeddings that maintain the original dimension and are then processed by a feed-forward network into box coordinates and class labels. By utilizing self-attention and encoder-decoder attention mechanisms, the model globally reasons about all the objects at once, considering their pairwise relations while leveraging the entire image as context.
@@ IV-B Predicting Object Relations as Graph Edges
Figure 3: Schematic illustration of our D3G. We can identify three main rows in this figure. The top row describes the transformer architecture composed of encoder and decoder. The former takes as input scene features and provides as output updated representations which are fed to the decoder together with a fixed number of object queries. The query features obtained from the decoder are filtered and matched to the ground truth objects in the scene to detect them during training. The middle row shows how the same object query representations define the edge features via pairwise concatenation and mappings, before being refined by several graph transformer layers. Finally, a binary classifier predicts from the edge features whether two objects in the scene are connected by a spatial dependency relation. The bottom row zooms inside a layer of the graph transformer assuming only one head for visual clarity.
@@
$A_{l,s}=\text{softmax}(\text{sum}_{z}(\hat{A}_{l,s}))~{}.$ (2)

### v2

@@ Abstract
Interacting with real-world cluttered scenes poses several challenges to robotic agents that need to understand complex spatial dependencies among the observed objects to determine optimal pick sequences or efficient object retrieval strategies. Existing solutions typically manage simplified scenarios and focus on predicting pairwise object relationships following an initial object detection phase, but often overlook the global context or struggle with handling redundant and missing object relations. In this work, we present a modern take on visual relational reasoning for grasp planning. We introduce D3GD, a novel testbed that includes bin picking scenes with up to 35 objects from 97 distinct categories. Additionally, we propose D3G, a new end-to-end transformer-based dependency graph generation model that simultaneously detects objects and produces an adjacency matrix representing their spatial relationships. Recognizing the limitations of standard metrics, we employ the Average Precision of Relationships for the first time to evaluate model performance, conducting an extensive experimental benchmark. The obtained results establish our approach as the new state-of-the-art for this task, laying the foundation for future research in robotic manipulation. We publicly release the code and dataset at https://paolotron.github.io/d3g.github.io.
@@ I Introduction
Previous research has framed this problem as visual manipulation relationship reasoning [1, 2, 3], highlighting as a primary challenge the inference of relationship among objects, which can be categorized as parent, child, or none. In this context, objects are nodes in a dependency graph, where parent nodes should be grasped after child nodes and this information is encoded in the edges. Knowing the complete dependency graph enables the calculation of the shortest path to any object and facilitates planning the most efficient picking sequence to access it.
@@
Existing algorithms address this task by decoupling object localization from relationship reasoning. They first identify the objects through a detector and then encode their relationship via a handcrafted post-hoc procedure known as object pair pooling [1, 2, 3]. This technique crops the learned feature map based on the union area of two objects to encode their relationship. However, in highly cluttered scenes, the cropping region may disproportionately represent one object over the other or include irrelevant objects. As a result, the generated representation is ambiguous and leads to inconsistencies in the relationship models, particularly in addressing issues such as redundant or missing edges [4, 1, 2].
While these shortcomings may not be immediately apparent in simplified scenarios involving a small number of objects (e.g. 5 objects as in [1]) or moderate setups (e.g. 19 objects in [5]) on a tabletop, they become increasingly evident as the task complexity scales. A bin filled with a large number of objects provides a compelling test case, as the bin’s boundaries facilitate object stacking, resulting in dense and intertwined layouts that correspond to highly complex dependency graphs. Moreover, the evaluation metrics commonly used to assess visual manipulation relationship reasoning methods are derived from detection literature and are influenced by an implicit object scoring threshold. This threshold is dataset-specific, challenging to calibrate, and may hinder the generalization of the evaluation process.
@@
(1) We introduce a new testbed based on the MetaGraspNetV2 dataset [6] that presents synthetic and real bin picking scenes annotated with dependency graphs. We manually organized the scenes in subsets of growing complexity by focusing on the number of nodes, connected subgraphs (clutter width), and shortest paths from leaf to root (clutter depth). We name it D3GD: Dependency Graph Generation for Grasp planning Dataset, and contains scenes with up to 35 objects from 97 different categories.
(2) We design a new end-to-end dependency graph generation model. We dub it D3G. Its objective is to jointly detect the objects and predict the adjacency matrix describing their spatial relationship. We formalize the task as an object set prediction problem over the entire image: leveraging global context, the model chooses where each object query should focus and a graph transformer ensures that relationship features attend to the relevant object queries. This enables simultaneous learning of object locations, appearances, and representation for their connecting edges.
(3) We assess the entire pipeline with a tailored metric. We use for the first time the Average Precision of Relationships [7] to evaluate the generated dependency graph. This metric calculates precision-recall trade-offs over predicted triplets defined by two objects and their relation. It provides a threshold-free evaluation score that extends beyond conventional object detection metrics. With a thorough experimental analysis we show how our approach outperforms several state-of-the art competitors in challenging experimental settings. Overall our work presents a novel benchmark suite that will be open-sourced to the community paving the way for future application-oriented research in robotic manipulation.
@@ II Related Works
Object Detection is an extensively studied task that aims to localize objects in a scene and recognize their category label. Detection approaches evolved from feature matching and part modeling [8], through two stages [9, 10] and one stage deep methods [11], getting more recently to transformer-based techniques. The latter allows for global computation over a whole image and set prediction, leveraging attention between multiple objects as in the case of DETR [12]. Deformable attention has proven to enhance such approaches [13]. Further enhancements have been obtained through query-denoising [14] and dynamic anchors [15, 16].
@@
Graph Learning effectively encodes complex interactions in structured data. Message passing based networks have been designed to leverage the sparsity and connectivity of graph structures [17], efficiently propagating information and capturing dependencies between nodes [18]. With the advent of self-attention based architectures [19], several efforts have been made to merge the graph neural network paradigm and the transformer one, obtaining competitive performance on graph benchmarks [20, 21, 22, 23].
@@
Relationship reasoning encompasses tasks where it is important to model geometric (e.g. spatial positioning) or semantic relations (e.g. subject-object interactions) from data. Large datasets such as Visual Genome [24] and Open Images [7] paved the way for perception studies mainly dedicated to image captioning, with first solutions building on two-stage object detectors [25], and the latest ones based on transformers [26] [27]. Several research works in robotics have been dedicated to dependency graph generation that aims at mapping the relationships and sequential constraints among multiple objects in highly cluttered scenes to support decision-making processes and motion planning [28, 29].
Such a graph is used as an intermediate result [30, 31] or an auxiliary objective for enhancing grasp detection models [32, 33]. In [34] dependency graphs are exploited to decide grasping order for efficient bin emptying, [35] uses them to predict the likelihood of object stacks collapse, and [36] leverages them to plan the optimal path to retrieve specific objects in dense clutter.
@@
The first real-world visual manipulation relationship dataset (VMRD) was introduced in [1]. It is a collection of 5185 hand-annotated images depicting various configurations of objects in cluttered scenes each containing up to 5 object instances sampled from 31 categories. While valid, this dataset is limited in dimension and complexity. The authors also proposed a deep model with distinct object detection and relationship prediction stages. The same high-level solution was then adopted by following approaches that used fully connected conditional random fields [4] and graph neural networks [2, 3]. These works included analyses on a larger relational grasp dataset named REGRAD [5] automatically collected in a virtual environment [^1] Reproducible results on a dataset require knowledge of the pre-processing pipeline and access to baseline code and checkpoints. Unfortunately, most prior work in visual manipulation relationship reasoning lacks these resources. Specifically, in REGRAD, the detailed analysis uncovered issues such as missing boxes, one-pixel masks, and inconsistent object classes. The lack of public code further impedes result replication. We overcome these limitations with our benchmark, re-implementing the competitor methods and releasing the code for fair comparison.. It comprises images annotated with manipulation relationships between objects in dense clutter and includes multi-view perspectives [37]. Nonetheless, existing testbeds tend to focus on relatively simple scenarios and often lack comprehensive information on the experimental settings, making previous results difficult to reproduce. Lately with the release of the NVIDIA Isaac simulator [38] a new generation of highly realistic and accurately annotated datasets have been published [39, 6, 40], which provide the opportunity to scale up relational reasoning related tasks for robotics applications.
@@ III Dataset
TABLE I: Difficulty levels of our D3GD dataset. Each row describes the concurrent conditions that define a level set.
Set # Objects Relative Clut. # Samples
Clut. Width Depth
@@
TABLE II: Overview of datasets for visual manipulation relationship reasoning, highlighting D3GD as the most challenging
Dataset Data Avg. Avg. Clutter Clutter
Type #Rel #Objs Depth Width
VMRD Train [1] Real 2.39 3.45 2.48 3.34
VMRD Test [1] Real 2.35 3.40 2.44 3.28
REGRAD Train [5] Synt 2.72 9.43 1.64 1.79
REGRAD Test [5] Synt 2.77 9.54 1.65 1.79
D3GD Train Synt 4.04 5.42 2.60 3.49
D3GD Test Easy Synt 2.89 4.89 2.41 2.84
D3GD Test Medium Synt 14.33 10.51 4.39 9.20
D3GD Test Hard Synt 30.40 16.05 4.46 15.05
D3GD Test Real Real 2.58 3.70 1.09 2.23
We define our experimental testbed D3GD to get closer to real-world conditions where robots are challenged with grasping objects from large containers. We build on MetaGraspNetV2 [6] which is composed of nearly 300K synthetic images and 2.3K real images depicting bin-picking scenes composed by up to 35 object instances sampled from 97 classes. The dataset comes with both 2D and 3D modalities as well as various types of grasp and labels.
For our use case, we consider only RGB images, the object bounding boxes, the grasping dependency graph, and, for the baseline models that use them, the instance segmentation annotations. We organize the dataset into different difficulty levels based on three key metrics: number of objects, clutter-depth, and clutter-width.
Figure 2: Data samples from the three main difficulty levels of D3GD. The objects in the bin on the left are identified by a number which is also used in the corresponding graph on the right. Here the arrows indicate a relationship dependency: A $\rightarrow$ B entails that A can’t be moved without first moving B. In a downstream grasp task, the objects that should be touched first are those that don’t have any outward relationships. For example, in the medium level, the green box (2) and brown box (6) can be safely grasped without causing object collapses that would alter the scene configuration.
To consider the scene density and the challenges provided by object occlusion we measure how many objects (nodes) are part of a connected subgraph within the whole dependency graph. We measure the clutter-depth as the longest shortest path from a graph leaf (the first object to be picked) to the root (the last one). This means that we count the minimum number of objects that we need to remove before reaching an object at the bottom of the bin. As the dependency graph of a scene may have multiple leaves and roots, we annotate for each scene the longest of these paths which is a good indicator of how many layers of clutter the bin contains. Note that we are interpreting the dependency graph as a Directed Acyclic Graph (DAG) and we removed all the not-fitting scene samples (only $\sim$60) that, upon visual inspection, were clearly identified as errors produced during the synthetic dataset generation phase. We also measure the clutter-width as the largest connected component: it yields a measure of how cohesive the clutter is, as seemingly complex scenes with many objects could become trivial when they are far apart and don’t have a relationship between them.
Overall, our D3GD is organized in difficulty sets by choosing manual thresholds upon visual inspection of the samples. In Tab. I we present the relative clutter width normalized by the number of objects in the scene. For our study, we disregard the trivial set and focus on the easy, medium, and hard ones (see Fig. 2). We split the dataset into train, eval, and test scenes with an 80:10:10 proportion.
@@
To frame our D3GD in the existing literature we present a comparison with the VMRD [1] and REGRAD [5] datasets on the basis of the metrics used to define the difficulty levels of our testbed. Tab. II shows how D3GD achieves complexities that weren’t explored before. In particular, even if the number of objects of REGRAD [5] is relatively high, the complexity metrics are all fairly low. This can be attributed to its simulated tabletop setting, where cluttered piles easily fall apart leading to sparsely arranged objects.
In contrast, our setting is characterized by denser, more intricate graphs, driven by the physical constraints of the bin environment.
@@ IV Method
In this section, we present our new visual relationship reasoning model, D3G, which is formalized as a transformer-based graph generation approach. It jointly predicts the graph nodes corresponding to objects in the scene, and the edges that encode their spatial relationships. Our end-to-end learning strategy avoids the drawbacks of two-stage pipelines [1, 2, 3] and allows the model to determine which are the most relevant information from the scene to solve the task.
We designed the architecture of D3G as a tailored integration of detection [12] and graph [21] transformers. The detection component extracts features through a combination of self and cross attention mechanisms, encoding them into a pre-defined number of object queries, set to exceed the maximum number of objects in the scene. These queries are pair-wise combined to represent object relationships. While incorporating an additional self-attention module to improve the relationship features via cross-talk could be beneficial, it becomes computationally prohibitive. Thus, we opted for a graph transformer that refines the relationship features iteratively, and we leveraged the fixed number of object queries to write its tensor operations into a dense format and obtain an efficient model.
Finally, the object queries are matched to the detected objects: the optimal assignment procedure allows us to evaluate the node loss as in standard detection [12], but we also extended its application to calculate the relationship loss. Specifically, the edge prediction matrix, initially generated through binary classification of all query pairs to determine relationship existence, is refined using a selection matrix defined by the optimal assignment.
Fig. 3 illustrates the proposed approach and helps to visualize the model flow explained in the following.
@@ IV-A Detecting Objects as Graph Nodes
Given an input image of dimension $3\times H_{0}\times W_{0}$, we perform object detection as a direct set prediction problem by following DETR [12]. A Convolution-based backbone extracts a compact feature representation of dimensionality $C\times H\times W$ (where $C=2048$, $H={H_{0}}/{32}$, $W={W_{0}}/{32}$) which is rescaled via a $1\times 1$ convolution to reduce the channel dimensionality. The resulting $d\times H\times W$ tensor is interpreted as a list of vectors ($d\times HW$). These are fed to a transformer encoder [19] together with their spatial positional encodings and the obtained output is passed to the transformer decoder which also receives $N$ object queries (tensor of dim. $N\times F_{0}$, initially set to zero). The latter is transformed by the decoder in new embeddings that maintain the original dimension and are then processed by a feed-forward network into box coordinates and class labels. By utilizing self-attention and encoder-decoder attention mechanisms, the model globally reasons about all the objects at once, considering their pairwise relations while leveraging the entire image as context.
@@ IV-B Predicting Object Relations as Graph Edges
Figure 3: Illustration of our D3G. The top row describes the transformer architecture composed of encoder and decoder. The former takes as input scene features and provides as output updated representations which are fed to the decoder together with a fixed number of object queries. The query features obtained from the decoder are filtered and matched to the ground truth objects in the scene to detect them during training. The middle row shows how the same object query representations define the edge features via pairwise concatenation and mappings, before being refined by several graph transformer layers. Finally, a binary classifier predicts from the edge features whether two objects in the scene are connected by a spatial dependency relation. The bottom row zooms inside a layer of the graph transformer assuming only one head for visual clarity.
@@
$A_{l,s}=\text{softmax}(\text{sum}_{z}(\hat{A}_{l,s}))~{}.\vspace{-1.5mm}$ (2)
We remark that operations analogous to those in Eq. (1) and (2) are managed in [21] by costly message passing layers. Instead, we are dealing with dense matrix operations performed via matrix multiplications which are easily parallelizable. This is possible because in our formulation we have a fixed set of queries as nodes in a fully connected graph.
@@
Figure 4: Relationship loss calculation procedure. The edge prediction matrix $e_{pred}$ contains a score for each pair of object queries. Only some of the N=5 queries match the real P=3 objects. The optimal bipartite matching $\hat{\sigma}=[5,3,1,2,4]$ identifies the correspondence between the predictions $\hat{y}=[3,0,2,0,1]$ and the ground truth objects $y=[1,2,3,0,0]$, which are in order Toy Part, Spray, and Soap. Here $T$ has dimensions $5\times 3$ and has value 1 in the cells that identify a match: e.g. cell (5,1) will contain a 1 because the object with ID 1 has been matched to the 5-th prediction. The matrix obtained by multiplying $T^{\top}\cdot e_{pred}\cdot T$ is directly comparable with the GT matrix via a simple cross-entropy loss.